OpenStudy (anonymous):

find the equation of the line joining the origin to the point of intersection 4x+3y=8 and x+y=1

OpenStudy (anonymous):

To find the point of intersection of the two lines we need to solve them and when we solve we get x = 8 and y = -7 so the point is (8,-7) then use the formula y - y1 / x - x1 = y2 - y1 / x2 - x1 so we get y - 0 / x - 0 = -7 - 0 / 8 - 0 which means y / x = -7 / 8 so 8y = -7x so the equation will be 8y + 7x = 0

OpenStudy (anonymous):

i want 2 confirm , did i write the solution correct or not

OpenStudy (anonymous):

plse help me friends

OpenStudy (anonymous):

Solve 4x+3y=8 and x+y=1 And Find x and y First

OpenStudy (anonymous):

i have already solve the question .plse check if i have done correct or not